The physical plant at the main campus of a large state university recieves daily requests to replace fluorescent lightbulbs. The distribution of the number of daily requests is bell-shaped and has a mean of 43 and a standard deviation of 4. Using the empirical rule, what is the approximate percentage of lightbulb replacement requests numbering between 43 and 47?
Answer #1

Solution :

Given that,

\mu = 43

\sigma = 4

P(43< x < 47) = P[(43 - 43) / 4< (x - \mu ) / \sigma \bar x < (47 - 43) /4 )]

= P(0 < Z <1 )

= P(Z < 1) - P(Z < 0)

using empirical rule

68% =  68%/2=34%

= P(Z < 1) - P(Z < 0)

=34% - 0%

=34%